OpenStudy (anonymous):

A person scores f(x) = 90/1 + 4e^-0.4t points on a test after t hours of studying. What does the person score without studying at all? Compute f’(0) and estimate how many points 1 hours of studying will add to the score.

OpenStudy (anonymous):

This is what I did but my teacher said it was wrong she said ( Find f' and f'(0).) F(x) = 90/ 1 + 4e^-0.4t F(0) = 90/ 1 + 4e^-0.4(0) = 18 F(1) = 90/1 + 4e^-0.4(1) = 24.45 So comparing the scote at f(1) and f(0) where at f(1) studying for 1 hour and f(0) studying none. F(1) - f(0)/24.45 – 18 = 6.45 points

OpenStudy (anonymous):

Should this have parenteses or not? Is this F(x) = 90/ (1 + 4e^-0.4t )?

OpenStudy (anonymous):

yes prenthesis

OpenStudy (anonymous):

like this \[\frac{ 9 }{ 1 + 4e ^{-0.4t}}\]

OpenStudy (anonymous):

Your teacher specifically said to differentiate?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

she said to Find f' and f'(0)

OpenStudy (anonymous):

@EulersEquation so do you know how?

OpenStudy (anonymous):

@Numb3r1 can u help?

OpenStudy (anonymous):

Yes. I need to review this. First glance is to use the quotient rule. Do you know how to use the quotient rule? Someone else may want to take a look. I am in the middle of helping someone else. When I finish I'll get on it.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Ha! I am terrible at arithmetic. On first run I get a negative number for f'(0). You cannot make a negative score on a test! Unless the teacher doesn't like you. I am rusty on this. I'll work on it some more. In the meantime, bump the problem and hopefully someone else will tackle this.

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

@waterineyes

OpenStudy (anonymous):

Just want to take derivative of that??

OpenStudy (anonymous):

yes Find f' and f'(0)

OpenStudy (anonymous):

It will come out to be negative..

OpenStudy (anonymous):

what will come out? so what i did above is wrong or?

OpenStudy (anonymous):

Wait, let we do it first..

OpenStudy (anonymous):

\[\frac{ 90 }{ 1 + 4e ^{-0.4t}} = \frac{ -(90)(-1.6 \cdot e^{-0.4t}) }{ (1 + 4e ^{-0.4t})^2}\]

OpenStudy (anonymous):

It will become positive now.. :)

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

Yeah. I don't see why it has to be differentiated -- not according to the question.

OpenStudy (anonymous):

Am I right in this?

OpenStudy (anonymous):

Ah! I forgot to square the denominator!

OpenStudy (anonymous):

That looks good.

OpenStudy (anonymous):

okay thanks!

OpenStudy (anonymous):

So you can now find f'(0) by putting 0 as value of t..

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

@onegirl can you go for that?

OpenStudy (anonymous):

just plug in 0 and solve right?

OpenStudy (anonymous):

Yep, just simple calculations with calculator..

OpenStudy (anonymous):

okay yes i can

OpenStudy (anonymous):

Good.. Keep it up..

OpenStudy (anonymous):

i got 0 after i solved